OpenStudy (gnrfan):

How can an expression written in either radical form or rational exponent form be rewritten to fit the other form?

OpenStudy (amistre64):

by recalling the rule for it i spose

OpenStudy (amistre64):

\[\Large \sqrt[k]{b^n}=b^{n/k}\]

OpenStudy (gnrfan):

k, I have a DBA with my teacher and it said be prepared for these types of questions

OpenStudy (gnrfan):

can you help me with another

OpenStudy (amistre64):

maybe .. depends on what the other is

OpenStudy (gnrfan):

How can the properties of rational exponents be applied to simplify expressions with radicals or rational exponents?

OpenStudy (amistre64):

i cant say for sure what a proper response would be for that. How can the properties be applied? By applying them is all i can think of to say about it. I mean, the properties arent going to do anyone any good if they are just left sitting around collect dust.

OpenStudy (amistre64):

the properties are just an extention of doing basic math. they are 'shortcuts' in a way that allows us to do operations without having to redo all the middle work

OpenStudy (gnrfan):

So like "You have to apply them all to the equation to get the full use and complete the problem with them." would be fine

OpenStudy (amistre64):

im not sure what would be fine :)

OpenStudy (amistre64):

i can math, writing complete sentences for math is not my strong point

OpenStudy (gnrfan):

complete opposite for me. Another or no?

OpenStudy (amistre64):

i can try another, but unless its a mathing problem i might not be able to be much use

OpenStudy (gnrfan):

some help is better than none. What reasoning and explanations can be used when solving radical equations? How do extraneous solutions arise from radical equations?

OpenStudy (amistre64):

your material might have a better answer for these, but i would say: radical equations are just exponent equation in disguise, so using properties of exponents may be useful. extra solutions tend to arise because when we attempt to get rid of a radical, we are messing around with a restricted domain and altering it.

OpenStudy (amistre64):

for example: sqrt(2x-4) = x+8 we can square each side to get rid of the radical 2x-4 = x^2+16x+64 the solutions to the new problem are from a different domain than the original setup. so we have to test our solutions to see what fits in the original

OpenStudy (gnrfan):

k, thanks for ur help. i know these aren't easy

OpenStudy (amistre64):

youre welcome. the site is acting up for me so communicating is not going so smoothly on my end at least. good luck with it all :)